# Tell Me What You See: An Iterative Deep Learning Framework for Image Captioning

### v1

@@ 4 Conclusion
In this paper, we have detailed the iterative design and evaluation of five image captioning models, charting a course from a simple baseline to a competitive attention-based system. Our systematic approach demonstrates a clear path for model enhancement and yields several key insights. First, we show that within the classic CNN-LSTM paradigm, simply upgrading a visual backbone is not a guaranteed path to better performance; it must be paired with an architecture, such as an attention mechanism, that can effectively process the increased feature complexity. Second, we show that the point of lowest validation loss does not always correspond to peak performance on task-specific metrics like BLEU, necessitating a nuanced model selection process.
While our final model, Nexus, does not compete with the scale or performance of contemporary Transformer-based systems like BLIP, its final BLEU-4 score of 31.4 validates our architectural design by surpassing several foundational benchmarks. This work’s contribution is not a new state-of-the-art result, but rather a clear, educational, and practical guide to understanding the foundational trade-offs in vision-language architecture that precipitated the field’s shift toward modern, attention-centric designs.
Table 4: Final Benchmark Comparison on the MS COCO Validation Set. Scores are reported as percentages. Our Nexus model is compared against foundational models and modern Transformer-based systems to provide a clear context for its performance.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR
Foundational CNN-LSTM Models
Google NIC (Vinyals et al., 2015) [5] 66.6 46.1 32.9 24.6
Soft-Attention (Xu et al., 2015) [6] 70.7 49.2 34.4 24.3 23.9*
Up-Down (Anderson et al., 2018) [7] 79.8 36.3 27.7*
Nexus (Ours, Epoch 13) 71.7 55.0 41.8 31.4 47.4
Transformer-Based Models (for reference)
OSCAR (Li et al., 2020) [21] 36.5 33.5*
BLIP (Li et al., 2022) [22] 39.7 36.1*
*METEOR scores from different eras/packages may not be directly comparable. Our NLTK-based METEOR score is reported for completeness.
@@
While the Nexus model demonstrates strong performance, several areas remain open for exploration. As an independent researcher, this work was developed with limited computational resources, which motivated our focus on efficient CNN-LSTM architectures and guided our evaluation choices. Future work could investigate lightweight or quantized variants of the architecture for deployment in resource-constrained environments. Second, generalization to out-of-distribution or low-resource image domains warrants deeper investigation, potentially using domain adaptation or contrastive pre-training techniques.
@@ Appendix A Qualitative Examples and Visualizations
This appendix provides additional visual results to complement the quantitative analysis in the main paper. We showcase attention maps from our best model, Nexus, to illustrate how the model dynamically focuses on relevant image regions during caption generation. We also provide more side-by-side caption comparisons to highlight the performance gains achieved through our iterative development process.

### v2

@@ 5 Conclusion
In this paper, this paper have detailed the iterative design and evaluation of five image captioning models, charting a course from a simple baseline to a competitive attention-based system. The systematic approach demonstrates a clear path for model enhancement and yields several key insights. First, this paper show that within the classic CNN-LSTM paradigm, simply upgrading a visual backbone is not a guaranteed path to better performance; it must be paired with an architecture, such as an attention mechanism, that can effectively process the increased feature complexity. Second, this paper show that the point of lowest validation loss does not always correspond to peak performance on task-specific metrics like BLEU, necessitating a nuanced model selection process.
While the final model, Nexus, does not compete with the scale or performance of contemporary Transformer-based systems like BLIP, its final BLEU-4 score of 31.4 validates the architectural design by surpassing several foundational benchmarks. This work’s contribution is not a new state-of-the-art result, but rather a clear, educational, and practical guide to understanding the foundational trade-offs in vision-language architecture that precipitated the field’s shift toward modern, attention-centric designs.
@@
While the Nexus model demonstrates strong performance, several areas remain open for exploration. As an independent researcher, this work was developed with limited computational resources, which motivated the focus on efficient CNN-LSTM architectures and guided the evaluation choices. Future work could investigate lightweight or quantized variants of the architecture for deployment in resource-constrained environments. Second, generalization to out-of-distribution or low-resource image domains warrants deeper investigation, potentially using domain adaptation or contrastive pre-training techniques.
@@ Appendix A Qualitative Examples and Visualizations
This appendix provides additional visual results to complement the quantitative analysis in the main paper. This paper showcase attention maps from the best model, Nexus, to illustrate how the model dynamically focuses on relevant image regions during caption generation.